OpenStudy (anonymous):

If a bell chimes once at 1 o clock and twice at 2 oclock and so on, on a 24 hour clock how many times?

OpenStudy (abb0t):

24?

OpenStudy (anonymous):

if in a day means

OpenStudy (anonymous):

$$1+2+3+\dots+12=\dots$$

OpenStudy (anonymous):

110

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

156

OpenStudy (anonymous):

156or 300

OpenStudy (anonymous):

\((1+2+3+\dots+12)+(1+2+3+\dots+12)=(1+12)+(2+11)+\dots+(12+1)=?\)

OpenStudy (zzr0ck3r):

\[\sum_{i=1}^{n}i=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

156 or 300 which is correct

OpenStudy (zzr0ck3r):

\[\sum_{i=1}^{24}i=\frac{24(24+1)}{2}=300\]

OpenStudy (zzr0ck3r):

unless im reading the problem wrong...

OpenStudy (anonymous):

oh what the hell it's a 24 hour clock... @zzr0ck3r is right

OpenStudy (mathstudent55):

I agree with @zzr0ck3r